Label the points O(0, 0), A(-2, 4), and B(1, 1) on the graph
of f (x) x 2. Determine the images of these points on the transformed function. Label the images O',A' and B'

f(x)=1/2x^2+2
1
Expert's answer
2020-12-06T10:28:20-0500

Label the points O(0, 0), A(-2, 4), and B(1, 1) on the graph

of f (x) = x2




transformation function

f(x) = "\frac{1}{2}" x2 +2


f(x) = "\frac{x^2}{2}" + 2


f(0) = "\frac{0^2}{2}" + 2 = 2

O' (0,2)


f(-2) = "\frac{(-2^2)}{2}" + 2 = 4


A' ( -2,4)


f(1) = "\frac{1^2}{2}" + 2 = "\frac{5}{2}" = 2.5


B' (1,2.5)